# Unveiling Linguistic Regions in Large Language Models

### v1

@@ 3.4 Monolingual Family Region
3.5 Further Pre-training Optimization
In this section, we demonstrate that stabilizing the core linguistic regions identified in Section 3.2 mitigates the catastrophic forgetting (CF) issue (McCloskey and Cohen, 1989; Kemker et al., 2018) in LLMs, while preserving performance comparable to full-scale fine-tuning in target language proficiency. Our experimental setup involves further pre-training LLaMA-2-7B on 100,000 Arabic sentences, with a batch size of $256$, a maximum token count of $512$, and learning rates ($lr$) and $5e{-5}$ or $5e{-6}$, employing perplexity (PPL) as the evaluation criterion.
Full-scale Model Fine-tuning
Traditional full-scale fine-tuning, when increasing the learning rate, enhances learning in the target language but aggravates forgetting in non-target languages. To counteract this forgetting, it is often essential to incorporate a portion of data from these other languages. As depicted on the left side of Figure 5 in line Blue, increasing $lr$ from $5e{-6}$(dotted blue line) to $5e{-5}$ (solid blue line) under full-scale fine-tuning boosts the acquisition of the target language (Arabic), while accelerates the forgetting rate of the non-target languages (English and Chinese) simultaneously, shown in the middle and right side.
Freeze Core Regions Fine-tuning
We hypothesize that CF problem occurs due to the amplification of parameter adjustments when increasing the learning rate, which leads to significant shifts in the core linguistic region, adversely affecting language alignment. To mitigate this, we protect the core linguistic region and key dimensions by freezing the ‘Top 5%’ core language area for fine-tuning, as shown by the red line in Figure 5. At a $lr$ of $5e{-6}$ (dotted line), the difference between freezing fine-tuning and full-scale fine-tuning is minimal. However, when the $lr$ increases to $5e{-5}$ (solid line), freezing fine-tuning not only facilitates faster learning in the target language (achieving better performance in Arabic PPL: 3.557 vs. 3.566), but also significantly reduces the forgetting of non-target languages (showing improvements in English and Chinese PPL: 18.796 vs. 20.557 and 90.84 vs. 563.423 , respectively).
@@ 3.5 Further Pre-training Optimization
Table 5: LLaMA-2-7B perplexity on 11 languages with a Russian region removal. Here, ‘Arabic’ and ‘Persian’ are gray-filled while others are unfilled, ‘Top’ and ‘Bottom’ are deduplicated, and ‘Base’ is unchanged. Values with greater changes compared to the other regions’ removals are in bold.
@@ 4 Realated Work
Neuron Importance Estimating
Several works have empirically estimated neuron importance. An effective importance metric is to utilize parameter magnitude (Zhu and Gupta, 2018; Renda et al., 2020; Zafrir et al., 2021). However, such a simple approach may inadequately measure a weight’s impact on model output. Another importance metric is to estimate sensitivity of parameters, essentially approximates the change in loss when a parameter is zeroed out (Molchanov et al., 2019; Sanh et al., 2020; Liang et al., 2021; Sapkota and Bhattarai, 2023). These studies all focus on pruning unimportant parameters through parameter importance estimation. In this work, we select the most crucial parameters to unveil the core linguistic and monolingual regions.
@@
Multilingual language models exhibit significant zero-shot and few-shot cross-lingual transferability across diverse tasks (Pires et al., 2019; Wu and Dredze, 2019; Xu et al., 2023). Fine-tuned on one language enables model to obtain comparable capabilities in another language (Muennighoff et al., 2023; Ye et al., 2023), often displaying code-switching behavior in context generation (Khanuja et al., 2020; Zhao et al., 2024). While enhancements in cross-lingual generalization through parameter and information transfer learning (Üstün et al., 2020; Choenni et al., 2023), compulsory language alignment (Sherborne and Lapata, 2022; Shaham et al., 2024) and in-context learning techniques (Winata et al., 2021; Tanwar et al., 2023) have been effective, comprehensive understanding of the internal mechanisms enabling cross-linguistic alignment in large language models (LLMs) is still lacking.
@@
Researchers have investigated the mechanisms underlying strong cross-lingual performance. Prior works have shown that multilingual multilingual language models rely on a shared subword vocabulary and joint pre-training across multiple languages (Pires et al., 2019; Cahyawijaya et al., 2023; Wu and Dredze, 2019). However, new insights highlight these models’ capacity for learning universal semantic abstractions (Artetxe et al., 2020; Chi et al., 2020) and demonstrate that mBERT (Devlin et al., 2019) embeddings of similar words in similar sentences across languages are approximately aligned already (Cao et al., 2020; Conneau et al., 2020; Xu et al., 2022). Analysis from a hierarchical perspective reveals that classifiers linked to different BERT (Devlin et al., 2019) layers assess semantic features through varied probe tasks (Lin et al., 2019; Jawahar et al., 2019). In this work, we introduce a parameter partitioning perspective within LLMs, identifying core linguistic and monolingual regions, which underpin cross-lingual alignment and language-specific characteristics, respectively.
@@ Appendix A Languages in Evaluation Corpus
Appendix B Evaluation on 30 Languages
The regions are localized from six languages: Arabic, Spanish, Russian, Chinese, Korean, and Vietnamese.
@@ Appendix B Evaluation on 30 Languages
Appendix C Remove and Freeze Fine-tuning
In Table 6, we expand removing-freezing ratio to $5\%$ for LLaMA-2-7B fine-tuning, and also observe that removing the ‘Top 5%’ regions destroys the original cross-lingual generalization competence.
Testing Dataset (Language) # Training Samples (Chinese) Removal Ratio $=5\%$
Top & Freeze Bottom & Freeze Top & Unfreeze
Wechat (Chinese) 0K 204630.063 6.454 204630.063
2K 816.94 6.057 6.05
5K 302.806 6.039 6.054
10K 167.891 6.285 6.315
20K 57.355 6.525 6.56
50K 18.92 6.154 6.173
200K 8.8 5.582 5.584
Falcon (English) 0K 66877.711 14.027 66877.711
2K 248218.656 14.486 14.44
5K 476895.25 14.831 15.074
10K 467330.656 16.248 16.334
20K 131648.75 18.579 18.743
50K 25494.166 29.576 29.377
200K 6948.663 321.353 504.645
Table 6: Removing-freezing analysis at $5\%$ removal ratio in different regions of LLaMA-2-7B. ‘Top/Bottom’ denotes the removal region, while ‘Freeze/Unfreeze’ indicates whether the corresponding region is frozen after removal.
Appendix D Attn. Dimensional Removal Evaluation
Figure 6 (left) illustrates that the columns of the Attn.k/q/v matrices in the attention layer, as well as the rows of the Attn.o matrix, correspond to different attention head parameters. Conversely, the rows of the Attn.k/q/v matrices and the columns of the Attn.o matrix are closely associated with features in the representation space.
We remove the ‘Top’ dimensions in the attention layer, and the results is displayed in Tables 7 and 8. Table 7 reveals that removing the ‘Top’ dimensions continues to produce more detrimental effects than other dimensions. The visualizations in Figure 2 show that these dimensions are largely concentrated in a few attention heads, suggesting that some attention heads contribute more significantly to the model’s linguistic competence. Table 8 indicates that the removals under the second setting cause more damage than the first. Considering that, in the second setting, the ‘Top’ dimensions in the matrix directly interact with the corresponding features in the representational space, we can conjecture that these features are tightly linked with the model’s linguistic competence.
@@ Appendix C Remove and Freeze Fine-tuning
Appendix E Single Parameter Perturbation
In a Transformer block, each column in the Attn.o and the MLP.down matrix of the feed-forward layer can be considered as the input weights of a neuron. Thus, perturbing a column can be seen as disturbing the input weights of a neuron. Viewed from another angle, if we disturb the output activation value of this neuron, a similar effect should be observed. Within LLaMA, there is a specific module called RMSNorm, where each dimension is associated with a weight. Perturbations to these weights can be regarded as disturbances to the output activation values of the corresponding neurons (In Figure 6 (right), we visually demonstrate how RMSNorm affects a column of the Attn.o and the FFN.down matrix).
Appendix F Individual Language Family Region
Tables 9-13 demonstrate LLaMA-2-7B perplexity after removing Arabic, Spanish, Chinese, Korean, and Vietnamese regions, respectively. The region is obtained by removing the intersections with other languages’ respective regions from the 1% ‘Top/Bottom’ regions, selected from 10,000 or 100,000 sentences during further pre-training according to Equation 4.

### v2

@@ 3.4 Monolingual Region
In Appendix F, we visualize the distribution Attn.q matrix in ‘Arabic’ and ‘Vietnamese’ regions and discover minimal overlap between them.
@@
Table 5: LLaMA-2-7B perplexity on 11 languages with a Russian region removal. Here, ‘Russian’ and ‘Ukrainian’ are gray-filled while others are unfilled, ‘Top’ and ‘Bottom’ are deduplicated, and ‘Base’ is unchanged. Values with greater changes compared to the other regions’ removals are in bold.
Downstream Task
We conduct downstream tasks on MMLU (Hendrycks et al., 2021) and ArabicMMLU (Koto et al., 2024). The former is in English while the latter is in Arabic. Our experiments demonstrate that removing monolingual regions significantly impacts the model’s ability to perform downstream tasks in the targeted language. Specifically, our findings are as follows:
1) The average logits probability for option ‘A’ on the ArabicMMLU evaluation increases from 0.36 to 0.64. 2) For the evaluation where the correct answer is one of the options ‘B/C/D/E’ in ArabicMMLU, the model’s accuracy decreases dramatically from 25.6% to just 1.5% when the ‘Arabic’ region is removed. However, if the ‘Vietnamese’ region is removed, the model’s accuracy remains almost unaffected at 26.7%. 3) For the MMLU evaluation, the model’s accuracy is minimally impacted by the removal of monolingual regions. Compared to a baseline accuracy of 42.46%, the removals of the ‘Arabic’ and ‘Vietnamese’ regions result in accuracies of 39.27% and 39.68%, respectively.
Furthermore, using "There are 365 days in a year and 12" as a prompt for generation, we test outputs on the removal of the removing monolingual regions. More details are in Appendix F.
Figure 5: Model’s accuracy on MMLU and ArabicMMLU (filtered) test. Here, ‘filtered’ denotes removing questions whose correct answer is ‘A’.
3.5 Further Pre-training Optimization
In this section, we demonstrate that stabilizing the core linguistic regions identified in Section 3.2 mitigates the catastrophic forgetting (CF) issue (McCloskey and Cohen, 1989; Kemker et al., 2018) in LLMs, while maintaining learning proficiency comparable to full-scale fine-tuning in target language. Our experimental setup involves further pre-training LLaMA-2-7B on 100,000 Arabic sentences, with a batch size of $256$, a maximum token count of $512$, and learning rates ($lr$) and $5e{-5}$ or $5e{-6}$, employing perplexity (PPL) as the evaluation criterion.
Full-scale Model Fine-tuning
Traditional full-scale fine-tuning, when increasing the learning rate or the amount of corpus data, enhances learning in the target language but aggravates forgetting in non-target languages. As shown in Table 2, since LLaMA-2 is primarily trained on English corpora, conducting a second stage of pretraining solely on large-scale Chinese corpora can lead to the forgetting of English competence. Additionally, as depicted on the left side of Figure 6 in blue line, increasing $lr$ from $5e{-6}$ (dotted line) to $5e{-5}$ (solid line) under full-scale fine-tuning boosts the acquisition of the target language (Arabic), while simultaneously accelerates the forgetting rate of the non-target languages (English and Chinese), as shown in the middle and right side.
Freeze Core Regions Fine-tuning
We hypothesize that CF problem occurs due to the amplification of parameter adjustments when increasing the learning rate, which leads to significant shifts in the core linguistic region, adversely affecting language alignment. To mitigate this, we protect the core linguistic region and key dimensions by freezing the ‘Top 5%’ core language area for fine-tuning, as shown by the red line in Figure 6. At a $lr$ of $5e{-6}$ (dotted line), the difference between freezing fine-tuning and full-scale fine-tuning is minimal. However, when the $lr$ increases to $5e{-5}$ (solid line), freezing fine-tuning not only facilitates faster learning in the target language (preserving comparable performance in Arabic PPL: 3.557 vs. 3.566), but also significantly reduces the forgetting of non-target languages (showing improvements in English and Chinese PPL: 18.796 vs. 20.557 and 90.84 vs. 563.423, respectively).
@@ 4 Realated Work
Intrinsic Regions
Prior works aimed to extract a sub-network capable of executing specific downstream tasks (Frankle and Carbin, 2019; Zhang et al., 2022) or task-specific subspaces to limit parameter fine-tuning within it (Aghajanyan et al., 2021; Zhang et al., 2023). For parameter importance estimation, an effective metric is to use parameter magnitude (Zhu and Gupta, 2018; Renda et al., 2020; Zafrir et al., 2021). Another metric involves estimating the sensitivity of parameters (Molchanov et al., 2019; Sanh et al., 2020; Liang et al., 2021; Sapkota and Bhattarai, 2023). In this work, we employ the latter method to select the most crucial parameters to unveil linguistic regions. (Frankle and Carbin, 2019) extracted a complete sub-network for downstream tasks, as exemplified by the optimal scale of the lottery network at 21.2%. However, our findings indicate that the linguistic regions is a much smaller (1%) functional region.
@@
Multilingual language models exhibit significant zero-shot and few-shot cross-lingual transferability across diverse tasks (Pires et al., 2019; Xu et al., 2023). Fine-tuned on one language enables model to obtain comparable capabilities in another language (Muennighoff et al., 2023; Ye et al., 2023), often displaying code-switching behavior in context generation (Khanuja et al., 2020; Zhao et al., 2024). While enhancements in cross-lingual generalization through parameter and information transfer learning (Üstün et al., 2020; Choenni et al., 2023), compulsory language alignment (Sherborne and Lapata, 2022; Shaham et al., 2024) and in-context learning techniques (Winata et al., 2021; Tanwar et al., 2023) have been effective, a comprehensive understanding of the internal mechanisms enabling cross-linguistic alignment in LLMs is still lacking.
@@
Prior works have shown that multilingual LMs rely on a shared subword vocabulary and joint pre-training across multiple languages (Pires et al., 2019; Cahyawijaya et al., 2023; Wu and Dredze, 2019). However, new insights highlight these models’ capacity for learning universal semantic abstractions (Artetxe et al., 2020; Chi et al., 2020) and demonstrate that mBERT (Devlin et al., 2019) embeddings of similar words in similar sentences across languages are approximately aligned already (Cao et al., 2020; Conneau et al., 2020; Xu et al., 2022). Analysis from a hierarchical perspective reveals that classifiers linked to different BERT (Devlin et al., 2019) layers assess semantic features through varied probe tasks (Lin et al., 2019; Jawahar et al., 2019). In this work, we introduce a parameter partitioning perspective within LLMs, identifying core linguistic and monolingual regions, which underpin cross-lingual alignment and language-specific characteristics, respectively.
Outlier Dimensions
Multiple studies have identified outlier dimensions in Transformer-based LMs (Kovaleva et al., 2021; Dettmers et al., 2022). Researches have found that certain outlier dimensions in PLMs are highly sensitive to the fine-tuning of downstream tasks (Kovaleva et al., 2021; Puccetti et al., 2022). Furthermore, (Rudman et al., 2023; Puccetti et al., 2022) discovered that these outlier dimensions encode task-specific knowledge, and disabling these dimensions significantly degrades model performance. (Luo et al., 2021) attribute positional embeddings to the emergence of outliers, while (Rajaee and Pilehvar, 2022) reported inconsistent results. Additionally, (Sun et al., 2024) found that these outlier dimensions exhibit significantly larger activation values than others in LLMs. Our findings in Section 3.3 demonstrate that beyond the outlier dimensions, other non-outlier dimensions within the core linguistic region also play an indispensable role in the model’s core linguistic competence.
@@ 5 Conclusion
Acknowledgements
The authors wish to thank the anonymous reviewers for their helpful comments. This work was partially funded by National Natural Science Foundation of China (No.62076069,62206057), Shanghai Rising-Star Program (23QA1400200), and Natural Science Foundation of Shanghai (23ZR1403500).
@@ Appendix A Languages in Evaluation Corpus
Appendix B Core Linguistic Region
The regions are localized from six languages: Arabic, Spanish, Russian, Chinese, Korean, and Vietnamese, respectively. Our work does not alter the embedding layer, as we think it equates to a mapping of tokens, which does not involve modeling linguistic competence.
Region Visualization
In Figure 9, we present the distribution of the ‘Top’ 5% regions of the LLaMA-2-13B model on the Attn.o matrix. The results indicate that across various layers, the core linguistic region on Attn.o matrix is concentrated on different rows. This difference is observed among the 40 attention heads in the 13B model.
@@ Appendix B Core Linguistic Region
Appendix C Attention Dimensional Removal
Figure 7 (left) illustrates that the columns of the Attn.k/q/v matrices in the attention layer, as well as the rows of the Attn.o matrix, correspond to different attention head parameters. Conversely, the rows of the Attn.k/q/v matrices and the columns of the Attn.o matrix are closely associated with dimensional features in the representation space.
We remove the ‘Top’ dimensions in the attention layer, and the results is displayed in Tables 6 and 7. Table 6 reveals that removing the Attention layers’ ‘Top’ dimensions continues to produce more detrimental effects than other dimensions. The visualizations in Figure 2 show that these dimensions are largely concentrated in a few attention heads, suggesting that some attention heads contribute more significantly to the model’s linguistic competence. Table 7 indicates that the removals under the second setting cause more damage than the first. Considering that, in the second setting, the ‘Top’ dimensions in the matrix directly interact with the corresponding dimensional features in the representational space, we can conjecture that these features are tightly linked with the model’s linguistic competence.
@@ Appendix C Attention Dimensional Removal
Appendix D Single Parameter Perturbation
In a Transformer block, each column in the Attn.o and the MLP.down matrix of the feed-forward layer can be considered as the input weights of a neuron. Thus, perturbing a column can be seen as disturbing the input weights of a neuron. Viewed from another angle, if we disturb the output activation value of this neuron, a similar effect should be observed. Within LLaMA, there is a specific module called RMSNorm, where each dimension is associated with a weight. Perturbations to these weights can be regarded as disturbances to the output activation values of the corresponding neurons. In Figure 7 (right), we visually demonstrate how RMSNorm affects a column of the Attn.o and the FFN.down matrix.
Perturbation Parameter Perplexity
- - 5.865
Reset 1 L1-N2100 83224.078
Reset 1 L1-N2800 5.860
Reset 1 L1-N4200 5.858
Mul 10 L1-N2100 4363.462
Mul 10 L1-N2800 5.859
Mul 10 L1-N4200 5.864
Appendix E Ablation Study
Tables 9 illustrate the perplexity of LLaMA-2-7B after removing core regions with and without outlier dimensions, respectively.
The ablation experiments reveal that different methods of disruption and varying model sizes exhibit different rates of PPL collapse:
1) Removing according to Attention.Head (attn.k/q/v.col + attn.o.row) results in a slower collapse than according to Dimensional Features (attn.k/q/v.row + attn.o.col). 2) The 13B model shows a slower rate of collapse. 3) The abnormal dimension is mainly concentrated in the FFN layer of the “core linguistic region”. If preserving outlier dimension, the speed of PPL collapse by removing FFN layers decreases most obviously, while Attention.Head is almost unaffected.
Removal Region $N_{d}$ LLaMA-2-7B Top(100K)
w/ outlier d w/o outlier d
Attn.o(row) Attn.k/q/v(column) FFN.down(coulumn) 1 848.326 27.265
3 72594.445 57308.313
5 48001.992 44730.059
10 62759.516 73425.438
Attn.o(row) Attn.k/q/v(column) 1 9.731 9.732
3 25.82 25.822
5 62.794 23.296
10 875.016 860.645
Attn.o(column) Attn.k/q/v(row) 1 167.804 9.586
3 68554.1 136.318
5 4259.861 688.476
10 68170.25 431317.863
FFN.up/gate(row) FFN.down(column) 1 20.039 6.727
3 74905.046 7.672
5 114725.578 9.946
10 239015.812 16.913
Appendix F Monolingual Region
Region Visualization
In Figure 10, we present the distribution of the Attn.q matrix for ‘Arabic’ and ‘Vietnamese’ in 4 different layers. The results reveal that across various layers, the two monolingual regions are concentrated in different columns of the matrix.
Region Removal
Tables 10-14 demonstrate LLaMA-2-7B perplexity after removing Arabic, Spanish, Chinese, Korean, and Vietnamese regions, respectively. The region is obtained by removing the intersections with other languages’ respective regions from the 1% ‘Top/Bottom’ regions, selected from 10,000 or 100,000 sentences during further pre-training according to Equation 4.
Case Study
In Figure 8, we use the prompt "There are 365 days in a year and 12" to test the model’s output in English, Arabic, and Chinese, respectively. The results indicate that removing the monolingual regions causes the model to lose the relative language competence, leading it to generate repetitive, nonsensical responses rather than correct answers like "12 months in a year".
Figure 8: Model’s generation with monolingual regions removed. Here, we use "There are 365 days in a year and 12" as prompt, and translate it into Arabic and Chinese to evaluate model’s performance in three languages.
Figure 9: Visualization of the linguistic competence region (the ‘Top’ 5% region) on Attention.o matrix across 8 different layers. The scale from 0 to 1 (after normalization) represent the proportion of parameters within a $3\times 3$ vicinity that belong to the ‘Top’ region.
Figure 10: Visualization of the monolingual regions for ‘Arabic’ and ‘Vietnamese’ across 4 different layers of the Attn.q matrix. The scale from 0 to 1 (after normalization) represent the proportion of parameters within a $3\times 3$ vicinity that belong to the monolingual regions.